If it take Karen 2 hours to wax SUV and Clarann 3 hours how long would it take them to wax it together.

Answer by Boreal(15235)

in 1 hour
Karen does (1/2) and Clarann does (1/3)
In x hours (x/2) and (x/3)
Set those two, their sum, equal to 1
(x/2)+(x/3)=1
multiply everything by LCD of 6
3x+2x=6
5x=6
x=6/5=1.2
it would take them together 1.2 hours, for it would be 1.2/2 (60%) for K and 1.2/3 (40%) for C.